Find the fifth, ninth and hundredth term of the sequence.
-19, -22, -25, -28,
A(n) = -19 + (n - 1)(-3)
I can't get this question correct, apparently. D:

Answer by stanbon(75887)

Find the fifth, ninth and hundredth term of the sequence.
-19, -22, -25, -28,
A(n) = -19 + (n - 1)(-3)
------
A(5) = -19+ 4(-3) = -19-12 = -31
---
A(9) = -19 + 8(-3) = -19-24 = -42
---
A(100) = -19 + 99(-3) = -19-297 = -316
-----------------------
Cheers,
Stan H.
----------